A simple pendulum consists
of a small object of mass 3.0 kg
hanging at the end of a 2.0-m-long
light string that is connected to a pivot point. (a) Calculate the magnitude of the torque (due to the force of gravity) about this pivot point when the string makes a 5.0° angle with the vertical. (b) Does the torque increase or decrease as the angle increases? Explain.
1
Expert's answer
2020-04-06T09:04:16-0400


a)Moment is equal

"M_1=m g \cdot {ON_1}=m g \cdot L \sin{\alpha_1 }=3 \cdot 9.81 \cdot 2 \cdot \sin{5^0 }=5.13 Nm"

b) When "\alpha_2 > \alpha_1" then "ON_2>ON_1" as "M_2= mg \cdot ON_2" then "M_2>M_1"

That is, with increasing angle, the moment increases